Cost of debt using both methods Currently, Warren Industries can sell 15-year, $1,000-par-value bonds paying annual interest at a 7% coupon rate. Because cur rent market rates for similar bonds are just under 7%, Warren can sell its bonds for $1,010 each; Warren will incur flotation costs of $30 per bond in this process. The firm is in the 40% tax bracket a. Find the net proceeds from sale of the bond, Nd. b. Show the cash flows from the firms point of view over the maturity of the bond. c. Calculate the before-tax and after-tax costs of debt d. Use the approximation formula to estimate the before-tax and after-tax costs of debt. e. Compare and contrast the costs of debt calculated in parts c and d. Which pg-2 approach do you prefer? Why?
Answer #1

a)

Selling Price $1010
Flotation Cost $30
Net Proceed from sale of the Bond $980

b) Below are the cash flows from the firm's point of view over the maturity of the bond

Time Cash Flow
0 $980
1-15 -70
15 -1000

c)

〉 Given [ユー高

d)

Before Tax After Tax
Annual Interest $70 (A)
Par Value $1000 (B)
Current Price of the bond $980 (C)
Duration 15 (D)
Tax rate 40% 40%
Yield to maturity(costs of debt)

=(A + ((B – C)/D))/((B + C)/2)

=(70 + ((1000 – 980)/15))/((1000 + 980)/2)

=(70 + 1.33)/(990) = 7.21%

= 7.21%*(1-0.4)

=4.326%